# Heterogeneous Mixture of Experts for Remote Sensing Image Super-Resolution

### v1

@@ I Introduction
High-resolution remote sensing imagery provides detailed representations of ground objects, serving critical functions in applications ranging from land-use monitoring [1] to post-disaster assessment [2] and ecological conservation. However, various constraints in imaging conditions and sensor hardware often result in the acquisition of low-resolution images [3], which significantly degrades their effectiveness in downstream applications. Super-resolution (SR) provides a non-physical solution to these resolution limitations by enabling the reconstruction of high-resolution (HR) images from low-resolution (LR) inputs [4].
Early super-resolution (SR) methods were primarily based on interpolation or optimization techniques[5, 6]. However, these approaches often struggle to model the complex nonlinear features present in remote sensing images. In recent years, deep learning-based SR methods have emerged as the dominant approach [7, 8, 9, 10], leveraging carefully designed deep neural networks to reconstruct images with fine structural details. Initially, convolutional neural networks (CNNs) were employed for SR reconstruction. Lideg et al. [11] introduced a CNN with residual connections, which was later modified by Lim et al [12]. Subsequently, attention-based methods gained widespread attention due to their superior long-range modeling capabilities. Zhang et al. proposed RCAN [13], which integrates convolutional operations with channel attention. Lei et al. developed a specialized self-attention mechanism to exploit self-similarity in remote sensing images. Chen et al. introduced HAT[9], which combines multiple attention mechanisms to extract more comprehensive features. Building on HAT, Xiao et al. [10] further incorporated token selection into the feature extraction process to reduce redundant information.
@@
However, current methods primarily rely on a single and fixed structure to process the intricate details of ground objects. This approach struggles to distinguish and reconstruct different types of ground objects effectively, often leading to suboptimal results. To address this limitation, we introduce a Mixture of Experts (MoE) model, where different experts are employed for upsampling.
@@
To better accommodate the complexity of ground objects in remote sensing images, we construct a set of heterogeneous experts. Specifically, these experts are grouped into multiple expert sets, where experts within the same group share the same structure, while those across groups have different architectures. This design allows for specialized upsampling tailored to different types of ground objects. Furthermore, to adapt to this heterogeneous expert design, we propose two key components. First, we introduce a Multi-level Feature Aggregation (MFA) strategy, which aggregates multi-level features from the backbone network to estimate expert activation probabilities, thereby guiding MoE routing. Second, we design a dual-routing mechanism, where an initial routing step selects the most suitable expert group, followed by a second routing step that determines the optimal expert within the selected group. Together, these components form the proposed Multi-level Feature Guided Heterogeneous Mixture of Experts (MFG-HMoE).
@@
2) To handle the complexity of remote sensing scenes, we design a set of heterogeneous experts and propose the MFA strategy along with the dual-routing mechanism to route ground object features to the most suitable expert for processing.
3) Our method outperforms contemporary approaches on the UCMerced [14] and AID [15] datasets, demonstrating its advanced capabilities.
@@ II-A Overview
We employ stacked Residual Hybrid Attention Groups (RHAG) [9] as the feature extraction backbone. To enhance the upsampling process, we introduce a Multi-Level Feature Aggregation (MFA) module that consolidates features from multiple levels to activate different experts. Within the upsampling process, we propose a Dual-Routing Heterogeneous Mixture of Experts (DR-HMoE) to enable adaptive pixel-level semantic expression, addressing the complex surface composition characteristics and diverse frequency patterns inherent in remote sensing imagery. Specifically, we design $N$ groups of upsampling experts, where experts within each group share identical structures while maintaining heterogeneity across groups. The expert activation follows a cascading process: first, the optimal expert group is selected based on MFA information, followed by the selection of the most suitable expert within that group using both MFA information and expert-specific cues. The detailed architecture is illustrated in Fig. 1.
@@ II-B Multi-level Feature Aggregation
where $L$ denotes the number of RHAG blocks in the backbone. The resulting $\mathbf{X}^{\text{agg}}$ serves as the MFA output and will be used as input to the router for activating different experts in the upsampling process.
@@ II-C Dual-Routing Heterogeneous Mixture of Experts
We propose the Dual-Routing Heterogeneous Mixture of Experts (DR-HMoE) to enhance the upsampling process in super-resolution networks. The DR-HMoE architecture comprises heterogeneous expert networks and their corresponding routers.
@@ II-C1 Heterogeneous Experts
where Conv2D denotes the convolutional operation and PS represents the pixel shuffle upsampling operation. To accommodate diverse types and scales of ground objects, experts across different groups utilize convolution kernels of varying sizes. While experts within each group share identical architectural structures, their parameters converge to distinct points in the parameter space during training. This heterogeneous architecture enables adaptive upsampling by providing specialized reconstruction functions for different pixels, thereby effectively handling various ground object patterns.
@@ II-C2 Dual-Routing Mechanism
where $j\in\{1,2,\cdots,M\}$ represents the $j$-th expert in the activated expert group $i$, $\mathbf{W}^{i}_{j}\in\mathbb{R}^{C\times M}$ represents the weights of the linear transformation for the $j$th expert in $i$-th group, and PE denotes the position encoding of the activated expert group. Based on the activation probabilities $p_{j}^{k}$, the top-K experts (denoted as $\Omega$) are selected to process the low-resolution features ($\mathbf{x}^{\text{feat}}_{k}$) according to:
$\mathbf{X}_{\text{out}}^{k}=\sum_{j\in\Omega}p_{j}^{k}(\mathbf{x}^{agg}_{k})E_{j}(\mathbf{x}^{\text{feat}}_{k})$ (7)
where $\mathbf{X}_{\text{out}}^{k}$ represents the output pixel value, and $\Omega$ denotes the indices of the selected experts. Through this approach, each pixel is adaptively routed to an expert group and its most suitable experts for upsampling and reconstruction.
@@
SSIM PSNR SSIM PSNR
EDSR [12] 0.9383 35.5104 0.7840 28.9274
@@
SSIM PSNR SSIM PSNR
EDSR [12] 0.9438 36.4926 0.8091 30.5434
@@
ID MFA Dual Routing Expert Num $1\times 1$ Experts $3\times 3$ Experts $5\times 5$ Experts PSNR SSIM
@@ III-A Datasets and Experimental Setup
For both datasets, we used the original images as high-resolution (HR) ground truth and generated the corresponding low-resolution (LR) images through bilinear interpolation at specific downsampling scales. We compared our approach with several state-of-the-art methods, including EDSR [12], RRDBNet [16], RCAN [13], SwinIR [17], HAT [9], HAUNet [18], and TTST [10].
@@ III-B Comparison with Other Methods
Tables I and II present the quantitative performance metrics for ×2 and ×4 SR tasks on both UCMerced and AID datasets. The results in Table I indicate that our method achieves substantial improvements over existing state-of-the-art approaches on the UCMerced dataset, particularly in the more challenging ×4 SR task. Furthermore, Table II shows that our method outperforms HAT in both SSIM and PSNR metrics without requiring modifications to the feature extraction network. Notably, these performance improvements exceed those achieved by TTST, despite TTST having a stronger feature extraction network.
@@ III-C Ablation Studies
Table III presents the ablation study results on the UCMerced dataset. The first row shows the baseline model with a single upsampling layer. The experimental results demonstrate several key findings: 1) Group 1 reveals the impact of MoE that incorporating 16 experts significantly improves the super-resolution performance. 2) Group 2 shows that adding MFA while maintaining 16 experts leads to further performance gains by effectively leveraging both deep and shallow features, thus maximizing the potential of the MoE model for super-resolution tasks. 3) Group 3 compares different numbers of experts, indicating that optimal performance requires an appropriate expert count, which depends on the dataset characteristics and feature distribution. 4) Group 4 demonstrates the crucial role of convolution sizes in upsampling, with $3\times 3$ kernels achieving optimal results. 5) Group 5 shows that introducing a dual routing mechanism further enhances super-resolution capability. While keeping the total number of experts constant, experiments with varying numbers and types of heterogeneous expert groups reveal that the combination of $1\times 1$ and $3\times 3$ heterogeneous experts yields the best performance.
@@ IV Conclusion
This paper presents a novel Multi-level Feature-Guided Heterogeneous Mixture of Experts (MFG-HMoE) model for remote sensing image super-resolution reconstruction. This model addresses the limitation of existing methods in distinguishing and reconstructing different ground object details. Our proposed framework employs multiple expert groups that maintain a balance between intra-group homogeneity and inter-group heterogeneity, facilitating specialized parameter activation for processing diverse ground details. Specifically, we introduce a Multi-level Feature Aggregation (MFA) module, which integrates deep and shallow features from the feature extraction network while providing crucial routing gate information for the Mixture of Experts (MoE). We propose a dual routing mechanism that optimizes upsampling expert assignment for individual pixels through a two-stage process. Experimental results on the UCMerced and AID datasets demonstrate that MFG-HMoE significantly surpasses state-of-the-art methods in super-resolution reconstruction accuracy. Through comprehensive ablation studies, we validate the effectiveness of each model component and establish optimal parameter configurations. This study advances the field of remote sensing image super-resolution reconstruction by introducing an innovative approach with promising implications for future research.

### v2

@@ I Introduction
High-resolution remote sensing imagery provides detailed representations of ground objects, serving critical functions in applications ranging from land-use monitoring [1, 2] to post-disaster assessment [3, 4] and ecological conservation [5, 6]. However, various constraints in imaging conditions and sensor hardware often result in the acquisition of low-resolution images [7], which significantly degrades their effectiveness in downstream applications. Super resolution (SR) provides a non-physical solution to these resolution limitations by enabling the reconstruction of high-resolution (HR) images from low-resolution (LR) inputs [8, 9].
Early super-resolution (SR) methods were primarily based on interpolation or optimization techniques[10]. However, these approaches often struggle to model the complex nonlinear features present in remote sensing images. In recent years, deep learning-based SR methods have emerged as the dominant approach, leveraging carefully designed deep neural networks to reconstruct images with fine structural details. Lei et al. first proposed LGCNet [8], a CNN-based remote sensing image SR method that reconstructs HR images by integrating local and global features. However, CNN-based methods suffer from limited receptive fields, making it difficult to model complex long-range dependencies in remote sensing images. To address this issue, a series of attention-based methods have been applied to remote sensing image SR, such as HSENet [11] and MSGFormer [12]. Furthermore, recent approaches have considered the characteristics of remote sensing images, including large variations in ground object scales and redundant details, further enhancing reconstruction performance through techniques such as token selection [13, 14], back projection [15], and saliency detection [16].
@@
However, current methods primarily rely on a single and fixed reconstruction structure to process the intricate details of ground objects. This approach struggles to distinguish and reconstruct different types of ground objects effectively, often leading to suboptimal results. The mixture-of-experts (MoE) [17, 18] model enables differentiated processing by assigning different inputs to specialized experts. Therefore, we introduce MoE into remote sensing image SR, employing a heterogeneous expert network in the upsampling reconstruction stage to achieve adaptive processing.
@@
To better accommodate the complexity of ground objects in remote sensing images, we construct a set of heterogeneous experts. Specifically, these experts are grouped into multiple expert sets, where experts within the same group share the same structure, while those across groups have different architectures. This design allows for specialized upsampling tailored to different types of ground objects. Furthermore, to adapt to this heterogeneous expert design, we propose two key components. First, we introduce a Multi-level Feature Aggregation (MFA) strategy, which aggregates multi-level features from the backbone network to estimate expert activation probabilities, thereby guiding MoE routing. Second, traditional single-step routing does not consider the inter-group heterogeneity of experts, which may lead to suboptimal results. To address this, we design a dual-routing mechanism, where an initial routing step selects the most suitable expert group, followed by a second routing step that determines the optimal expert within the selected group. Together, these components form the proposed Multi-level Feature Guided Heterogeneous Mixture of Experts (MFG-HMoE).
@@
2) To handle the complexity of remote sensing scenes, we design a set of heterogeneous experts and propose the MFA strategy with the dual-routing mechanism to route ground object features to the most suitable expert for processing.
3) Our method outperforms contemporary approaches on the UCMerced and AID datasets [19, 20, 21], demonstrating its advanced capabilities.
@@ II-A Overview
We employ stacked Residual Hybrid Attention Groups (RHAG) [22] as the feature extraction backbone. To enhance the upsampling process, we introduce a Multi-Level Feature Aggregation (MFA) module that consolidates features from multiple levels to activate different experts. Within the upsampling process, we propose a Dual-Routing Heterogeneous Mixture of Experts (DR-HMoE) to enable adaptive pixel-level semantic expression, addressing the complex surface composition characteristics and diverse frequency patterns inherent in remote sensing imagery. Specifically, we design $N$ groups of upsampling experts, where experts within each group share identical structures while maintaining heterogeneity across groups. The expert activation follows a cascading process: first, the optimal expert group is selected based on MFA information, followed by the selection of the most suitable expert within that group using both MFA and expert-specific cues. The detailed architecture is illustrated in Fig. 1.
@@ II-B Multi-level Feature Aggregation
where $L$ denotes the number of RHAG blocks and LN denotes the layer normalization. The resulting $\mathbf{X}^{\text{agg}}$ serves as the MFA output and will be used as input to the router for activating different experts in the upsampling process.
@@ II-C Dual-Routing Heterogeneous Mixture of Experts
We propose the Dual-Routing Heterogeneous Mixture of Experts (DR-HMoE) to enhance the upsampling process in SR networks. The DR-HMoE architecture comprises heterogeneous expert networks and their corresponding routers.
@@ II-C1 Heterogeneous Experts
where Conv2D denotes the convolutional operation and PS represents the pixel shuffle upsampling operation. To accommodate diverse types and scales of ground objects, experts across different groups utilize convolution kernels of varying sizes. While experts within each group share identical architectural structures, their parameters converge to distinct points in the parameter space during training. In another word, the inter-group heterogeneous experts provide different reconstruction scales, while the intra-group homogeneous experts offer multiple reconstruction patterns within the same scale. This architecture enables adaptive upsampling by providing specialized reconstruction functions for different pixels, thereby effectively handling various ground object patterns.
@@ II-C2 Dual-Routing Mechanism
where $j\in\{1,2,\cdots,M\}$ represents the $j$-th expert in the activated expert group $i$, $\mathbf{W}^{i}_{j}\in\mathbb{R}^{C\times M}$ represents the weights of the linear transformation for the $j$th expert in $i$-th group, and PE denotes the position encoding of the activated expert group. Based on the activation probabilities $p_{j}^{k}$, the top-$K$ experts in $i$-th group are selected to process the low-resolution features ($\mathbf{x}^{\text{feat}}_{k}$) according to:
$\mathbf{X}_{\text{out}}^{k}=\sum^{K}_{~{}j\in G_{i}}p_{i}^{k}(\mathbf{x}^{agg}_{k})p_{j}^{k}(\mathbf{x}^{agg}_{k})E_{ij}(\mathbf{x}^{\text{feat}}_{k})$ (7)
where $\mathbf{X}_{\text{out}}^{k}$ represents the output pixel value, and $G_{i}$ denotes the indices of the selected experts in $i$-th group. Through this approach, each pixel is adaptively routed to an expert group and its most suitable experts for upsampling.
@@
SSIM PSNR (dB) SSIM PSNR (dB)
EDSR-L [23] 0.9383 35.5104 0.7840 28.9274
@@
SSIM PSNR (dB) SSIM PSNR (dB)
EDSR-L [23] 0.9438 36.4926 0.8091 30.5434
@@
ID MFA Dual Routing Expert Num $1\times 1$ Experts $3\times 3$ Experts $5\times 5$ Experts SSIM PSNR (dB)
@@ III-A Datasets and Experimental Setup
For both datasets, we used the original images as high-resolution (HR) ground truth and generated the corresponding low-resolution (LR) images through bilinear interpolation at specific downsampling scales. We compared our approach with several state-of-the-art methods, including EDSR-L [23], RRDBNet [24], RCAN [25], SwinIR [26], HAT [22], HAUNet [27], SPT [16] and TTST [13].
@@ III-B Comparison with Other Methods
Tables I and II present the quantitative performance metrics for ×2 and ×4 SR tasks on both UCMerced and AID datasets. The results in Table I indicate that our method achieves substantial improvements over existing state-of-the-art approaches on the UCMerced dataset, particularly in the more challenging ×4 SR task. Furthermore, Table II shows that our method outperforms HAT in SSIM and PSNR metrics without requiring modifications to the feature extraction network. Notably, these performance improvements exceed those achieved by TTST, despite TTST having a stronger feature extraction network.
@@ III-C Ablation Studies
Table III presents the ablation study results on the UCMerced dataset. The first row shows the baseline model with a single upsampling layer. The experimental results demonstrate several key findings: 1) Group 1 reveals the impact of MoE that incorporating 16 experts significantly improves the super-resolution performance. 2) Group 2 shows that adding MFA while maintaining 16 experts leads to further performance gains by effectively leveraging both deep and shallow features, thus maximizing the potential of the MoE model for super-resolution tasks. 3) Group 3 compares different numbers of experts, indicating that setting the number of experts to 16 achieves the best reconstruction performance on the UCMerced dataset. 4) Group 4 demonstrates the crucial role of convolution sizes in upsampling, with $3\times 3$ kernels achieving optimal results. 5) Group 5 shows that introducing a dual routing mechanism further enhances super-resolution capability. While keeping the total number of experts constant, our experiments on the number of heterogeneous expert groups and their combinations reveal that the combination of $1\times 1$ and $3\times 3$ heterogeneous experts yields the best performance.
@@ IV Conclusion
This paper presents a novel Multi-level Feature-Guided Heterogeneous Mixture of Experts (MFG-HMoE) model for remote sensing image super-resolution reconstruction. This model addresses the limitation of existing methods in distinguishing and reconstructing different ground object details. Our proposed framework employs multiple expert groups that maintain a balance between intra-group homogeneity and inter-group heterogeneity, facilitating specialized parameter activation for processing diverse ground details. Specifically, we introduce a Multi-level Feature Aggregation (MFA) module, which integrates deep and shallow features from the feature extraction network while providing crucial routing gate information for the Mixture of Experts (MoE). We propose a dual routing mechanism that optimizes upsampling expert assignment for individual pixels through a two-stage process. Experimental results on the UCMerced and AID datasets demonstrate that MFG-HMoE significantly surpasses state-of-the-art methods in super-resolution reconstruction accuracy. Through comprehensive ablation studies, we validate the effectiveness of each model component and establish optimal parameter configurations.